Assignment 1 Wage Experience CollGrad 11 2 1 4 3 0 14 7 0 12 3 24 15 0 6 4 0 13. We will now change the scale of one of the variables. Continue to use the wage data above. • Convert wages in the table abplease explain all of 3 and 4 please
Answer #1
Wages Experience ColGrad
11 2 1
4 3 0
14 7 0
12 3 1
24 15 0
6 4 0

At the exchange rate 1 euro = 1.06125 dollars, we get the euro wages as -

Wages in dollar Wages in Euro = dollar wage / 1.06125 Experience ColGrad
11 10.365 2 1
4 3.769 3 0
14 13.192 7 0
12 11.307 3 1
24 22.615 15 0
6 5.654 4 0
Total = 69.902

Mean = 69.902 / 6 = 11.15 euros

This is the mean wage.

Now for variance -

Wages in Euro = dollar wage / 1.06125 (x-mean)^2
10.365 0.6162
3.769 54.48
13.192 4.17
11.307 0.025
22.615 131.45
5.654 30.21
Total = 69.902 Total = 220.95

Variance = 220.95 / 5 = 44.19

Now -

To prove -

COV(ax,by)=abCOV(x,y)

We have. -

COV(ax,by)=E[aX-E(aX)][bY-E(bY)]

\Rightarrow COV(ax,by)=E[aX-aE(X)][bY-bE(Y)]

\Rightarrow COV(ax,by)=aE[X-E(X)]b[Y-E(Y)]

\Rightarrow COV(ax,by)=(ab)E[X-E(X)][Y-E(Y)]

\Rightarrow COV(ax,by)=(ab)COV(x,y)

Hence, proved.